OpenStudy (anonymous):

New Algebra Question: 8.02] Determine the factors of a2b + 2a2 + 3b + 6. (b + 3)(a2 + 2) (b + 2)(a2 + 3) (ab + 2)(a + 3) (ab + 3)(a + 2)

OpenStudy (anonymous):

a2(b+2) + 3(b+2) => (a2+3)(b+2)

OpenStudy (anonymous):

Huh?

OpenStudy (anonymous):

So B?

OpenStudy (anonymous):

Huh what, those are the factors.

OpenStudy (anonymous):

B, yes.

OpenStudy (anonymous):

Oh, ok! :) How?

OpenStudy (anonymous):

This is 'factoring-by-grouping' \(\large a^2b + 2a^2 + 3b + 6\) Factor the first two terms separately from the last two terms. \(\large a^2b + 2a^2 = a^2(b+2)\) \(\large 3b + 6=3(b+2)\) (b+2) is a common factor of both.

OpenStudy (anonymous):

Ok, thx so much! :)